Fermi National Laboratory
 

Operation Summary for October 12, 2011 - October 14, 2011

 

Wednesday October 12

The day shift began with Operations monitoring protons in the Recycler (9.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Antiproton Source (Pbar) personnel began reverse proton studies around 9:28 AM. Operators reported at 10:47 AM that a Switchyard power supply (V204) was not ramping; a Power Supply person investigated and reported that the supply’s reference was in local. Operations established beam to the Meson FTBF experiment T-992 at 11:14 AM. Operators prepared MiniBooNE for an access at 11:57 AM; they came out of access at 2:44 PM. Operations prepared MTA for an access at 2:50 PM, but no one accessed the area. Pbar personnel ended their studies at 3:55 PM.

During the evening shift, at 6:24 PM, a MI Lambertson (I:LAM42) tripped off on excessive ripple; it reset. Operations disabled beam to T-992 at 10:05 PM.

Thursday October 13

The midnight shift began with Operations monitoring protons in the Recycler (10.8E10), and with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, at 12:53 AM, I:LAM42 tripped off on excessive ripple; it reset. The MI power supply interlock loop filter choke (MI4OL) tripped off; it reset. The Neutron Therapy Facility (NTF) began preparing at 7:30 AM for treating one patient later in the morning.

During the day shift, a Pbar person began a study period at 8:20 AM. Operations established beam to T-992 around 10:30 AM. At 1:47 PM, the MI beam permit tripped off when a kicker (KPS1C) prefired; it reset. Booster personnel conducted studies through out the shift. Pbar personnel completed their studies at 3:39 PM. At some point during the day, FESS personnel accessed MTA and replaced its faulty breaker trip unit.

During the evening shift, at 5:35 PM, the MiniBooNE beam permit tripped off when a power supply’s (E:HV860F) output drifted low; it reset. At 5:56 PM, the NuMI BPMs at MI-60 South stopped responding; a reboot brought them back online. Operations disabled beam to T-992 at 10:07 PM.

Friday October 14

The midnight shift began with Operations monitoring protons in the Recycler (5.9E10), and with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, at 12:39 AM, a Linac quadrupole power supply (L:QPSC6) tripped off; operators swapped controls cards with QPSC7 and had both supplies up and running within 10 minutes. Operations turned off beam to MiniBooNE at 5 AM and prepared the MI-12 area for a scheduled access later in the morning. Operators handed out keys for a MTA access at 7:13 AM. EE Support personnel reported at 7:24 AM that they had adjusted the MiniBooNE Horn power supply ground fault trip limit back to its nominal setting. NTF personnel began a tune-up at 7:30 AM in preparation for treating one patient later in the morning.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The plans for today are to turn on MiniBooNE and to keep everything running normally today and through the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/14/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory